# GIST: Gauge-Invariant Spectral Transformers for Scalable Graph Neural Operators

### 17 Mar 2026 v1

@@ Abstract
Adapting transformer positional encoding to meshes and graph-structured data presents significant computational challenges: exact spectral methods require cubic-complexity eigendecomposition and can inadvertently break gauge invariance through numerical solver artifacts, while efficient approximate methods sacrifice gauge symmetry by design. Both failure modes cause catastrophic generalization in inductive learning, where models trained with one set of numerical choices fail when encountering different spectral decompositions of similar graphs or discretizations of the same mesh. We propose GIST (Gauge-Invariant Spectral Transformers), a new graph transformer architecture that resolves this challenge by achieving end-to-end $\mathcal{O}(N)$ complexity through random projections while algorithmically preserving gauge invariance via inner-product-based attention on the projected embeddings. We prove GIST achieves discretization-invariant learning with bounded mismatch error, enabling parameter transfer across arbitrary mesh resolutions for neural operator applications. Empirically, GIST matches state-of-the-art on standard graph benchmarks (e.g., achieving 99.50% micro-F1 on PPI) while uniquely scaling to mesh-based Neural Operator benchmarks with up to 750K nodes, achieving state-of-the-art aerodynamic prediction on the challenging DrivAerNet and DrivAerNet++ datasets.
@@ 1 Introduction
Following their incredible success for processing sequential data in Natural Language Processing, Transformers (Vaswani2017) have been demonstrating a remarkable capacity for handling data of increasing structural complexity. Lee2019 have proposed a variant of the transformer block for permutation invariant data with their Set Transformer architecture; Dosovitskiy2021 have adapted the self-attention mechanism to 2D images with the very influential Vision Transformer architecture; and Bertasius2021 have extended transformers to video analysis with their Video Vision Transformer (ViViT), demonstrating how attention mechanisms capture both spatial and temporal dependencies across video frames. This progression from sequential text to increasingly structured data indicates a trajectory suggesting that Transformers are poised to tackle even more complex data structures, including irregular meshes and graphs.
Indeed, recent developments in adapting Transformers to graphs have shown promising results in capturing long-range dependencies that traditional Graph Neural Networks (GNNs) struggle with due to their reliance on localized message passing (Dwivedi2022; Zhu2023c). Unlike GNNs that aggregate information from nearest neighboring nodes by iterating through layers, Transformers can directly capture global relationships across the whole graph through self-attention, enabling them to reason about distant node interactions in a single layer.
@@
Consider a graph whose spectral embeddings are computed via random projection matrix $R$. A neural network trained on these embeddings $\{R\phi_{i}\}_{i}$ will inevitably learn features correlated with the specific choice of $R$. Consequently, when evaluated on different graphs or with a different random projection $R^{\prime}$ or different numerical eigensolver choices (sign flips, eigenspace ordering, handling eigenspace degeneracy), learned features become meaningless. This gauge dependence fundamentally undermines generalization, a critical failure mode for inductive graph learning where models must transfer to unseen structures.
Beyond graphs, gauge invariance is essential to generate neural operators with bounded discretization error. Physical problems (e.g., computational fluid dynamics, structural mechanics, shape analysis) are defined on continuous manifolds but discretized into computational meshes corresponding to graphs. Different mesh resolutions produce different graph Laplacians with different spectral decompositions, each involving arbitrary gauge choices (sign flips, eigenspace rotations, solver artifacts). Without gauge invariance, parameters trained on one discretization fail to transfer to another, and the attention kernels computed from spectral embeddings at different resolutions cannot be compared meaningfully. Gauge invariance ensures that the learned operator converges to the same continuum limit regardless of discretization, enabling provably bounded discretization mismatch error that vanishes as resolution increases.
@@
We propose Gauge-Invariant Spectral Transformers (GIST), which overcomes both barriers through a key insight: gauge invariance can be preserved by restricting attention to inner products between spectral embeddings, which remain invariant even under approximate spectral computations. We instantiate this framework using random projections for their computational efficiency and theoretical guarantees. For random projections, inner products between projected embeddings $\langle R\phi_{i},R\phi_{j}\rangle=\langle\phi_{i},(R^{T}R)\phi_{j}\rangle\approx\langle\phi_{i},\phi_{j}\rangle$ remain approximately invariant under gauge transformations. By restricting attention operations to only these inner products and combining with linear attention, we achieve end-to-end $\mathcal{O}(N)$ complexity while recovering gauge invariance algorithmically. Unlike prior approaches that address these barriers separately, GIST achieves gauge invariance and linear complexity simultaneously. Our contributions in short are:
@@
    identifying gauge invariance breaking as a fundamental limitation shared by both exact and approximate spectral methods, and characterizing when this breaks generalization and neural operator convergence;
@@
    proposing GIST, combining gauge-invariant spectral attention with multi-scale linear transformer blocks, achieving theoretical guarantees on complexity and invariance preservation, with competitive results on standard graph benchmarks (e.g. Cora, PubMed, PPI);
@@
    establishing GIST as a neural operator with provably bounded discretization mismatch error achieving state-of-the-art on large-scale mesh problems.
@@ 2 Related Works
Graphormer (Ying2021) introduces the idea of integrating structural encodings such as shortest path distances and centrality in Transformers. Similarly, Dwivedi2022 propose LSPE (Learnable Structural and Positional Encodings), an architecture that decouples structural and positional representations. Kreuzer2021 propose Spectral Attention Network (SAN), which introduces learned positional encodings from the full Laplacian spectrum. Park2022a develop Graph Relative Positional Encoding (GRPE), which extends relative positional encoding to graphs by considering features representing node-topology and node-edge interactions. Hierarchical Graph Transformer (Zhu2023c) addresses scalability to million-node graphs through graph hierarchies and coarsening techniques.
SpecFormer (bo2023specformer) and PolyFormer (Chen2025) are recent spectral Transformers that improve over SAN by leveraging approximate spectral bases or low-rank polynomial Laplacian filters to enhance scalability and accuracy on graph tasks.
Several recent architectures aim to bridge the gap between local and global reasoning via structural encodings and scalable attention. GraphGPS (rampasek2022GPS) adopts a hybrid paradigm that decouples local message passing from global attention while integrating spectral positional encodings to capture multi-scale interactions. However, unlike GIST, it lacks a mechanism to algorithmically preserve gauge invariance, leaving the model susceptible to the arbitrary numerical artifacts and sign flips inherent in spectral decompositions. While models like Exphormer (shirzad2023exphormer) attempt to address the resulting complexity by replacing full attention with sparse expander-based mechanisms, they still struggle with the high overhead of global coordinate systems. Similarly, tokenized approaches such as TokenGT (hamilton2017inductive) and NAGphormer (chen2023nagphormer) treat graphs as sets of tokens with [CLS]-style readouts, yet these models scale poorly due to memory-intensive tokenization and positional encoding costs that grow super-linearly with graph size. Consequently, these frameworks are rarely evaluated on large-scale inductive tasks like Elliptic or ogbn-arxiv, where the computational burden of maintaining structural encodings becomes prohibitive.
@@
Recently, neural operators have shown success in learning maps between continuous function spaces rather than fixed-dimensional vectors. Two properties are crucial here: (i) discretization invariance, i.e., a single set of parameters applies across discretizations (meshes, resolutions, and sampling locations) of the same underlying continuum problem; and (ii) global integration, i.e., the ability to represent nonlocal interactions via learned integral kernels, rather than being limited to finite-receptive-fields. Formally, a neural operator composes learned integral operators with pointwise nonlinearities, yielding universal approximation results for continuous nonlinear operators and implementations that share weights across resolutions. Our approach preserves these neural operator properties and improves scalability, allowing it to be applied to these cases (kovachki2023neuraloperator).
@@ 3.1 Preliminaries
The (normalized) graph Laplacian $\mathcal{L}=\mathds{1}-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$}} induces a natural metric via the resistance distance: $\Omega(i,j)=(e_{i}-e_{j})^{\top}\mathcal{L}^{\dagger}(e_{i}-e_{j})$, where $e_{i}$ is the $i$th standard basis vector and $\mathcal{L}^{\dagger}$ is the Moore-Penrose pseudoinverse (Klein1993).
The resistance distance can be expressed via Laplacian eigenmaps, which satisfy:
@@
with $\lambda_{k},u_{k}$ being the eigenvalues and eigenvectors of $\mathcal{L}$. These eigenmaps are natural positional encodings for graphs because their pairwise distances preserve the graph’s metric structure (Dwivedi2021). However, exact computation requires $\mathcal{O}(N^{3})$ eigendecomposition, prohibitive for large graphs.
@@
GIST’s gauge-invariant attention requires approximate spectral methods that preserve inner products between embeddings. We use FastRP (Chen2019e), which employs random projections $R\in\mathbb{R}^{r\times N}$ with $r=\mathcal{O}(\log(N)/\epsilon^{2})$ and $k$ power iterations to compute approximated eigenmaps $\tilde{\phi}_{i}=R\phi_{i}\in\mathbb{R}^{r}$ with $\mathcal{O}(N\log N)$ complexity.
By the Johnson-Lindenstrauss Lemma, while individual embeddings $\tilde{\phi}_{i}$ depend on the arbitrary choice of $R$, the inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle\approx\langle\phi_{i},\phi_{j}\rangle$ remain approximately invariant. This inner-product preservation is essential for gauge-invariant attention, ensuring attention weights do not depend on arbitrary numerical choices in the projection.
@@
While the approximate eigenmaps $\tilde{\phi}_{i}$ are efficient and preserve distances, as mentioned their gauge dependence is problematic: neural networks trained on these embeddings will learn features correlated with the specific projection matrix $R$. This creates a generalization failure in inductive settings where different graphs or different numerical solvers produce different projections.
Our approach addresses this by using approximate eigenmaps as positional encodings, but restricting the Transformer to operations that depend only on gauge-invariant quantities.
@@
The key insight is that while the projection matrix $R$ breaks gauge invariance, the inner products between projected embeddings remain approximately invariant: $(R\phi_{i})^{\top}(R\phi_{j})=\phi_{i}^{\top}(R^{\top}R)\phi_{j}\approx\phi_{i}^{\top}\phi_{j}$. By taking care that all operations depend on the embeddings only through these inner products we preserve gauge invariance by design while maintaining computational efficiency.
Figure 1: Gauge-Invariant Spectral Transformer. Left: Gauge-Invariant Spectral Self-Attention operates on graph positional embeddings $\tilde{\phi}$ as queries and keys, and node features $x$ as values. The output of the self-attention operation is then combined with $x$ through a residual connection. Limiting $\tilde{\phi}$ to queries and keys preserves gauge invariance across the self-attention block. Right: Gauge-Invariant Spectral Self-Attention is embedded in a Multi-Scale Gauge-Invariant Spectral Transformer Block which comprises 3 parallel branches inspired by EfficientViT.
@@
We now introduce our main contribution: Gauge-Invariant Spectral Transformer (GIST). The first ingredient of GIST is Gauge-Invariant Spectral Self-Attention, which operates on approximate spectral embeddings $\tilde{\phi}_{i}=R\phi_{i}\in\mathbb{R}^{r}$ but restricts attention computations to inner products between embeddings. The key observation is that while the embeddings themselves depend on the arbitrary projection matrix $R$, the inner products $(R\phi_{i})^{\top}(R\phi_{j})=\phi_{i}^{\top}(R^{\top}R)\phi_{j}\approx\phi_{i}^{\top}\phi_{j}$ are approximately gauge-invariant because $R^{\top}R\approx I$ by Johnson-Lindenstrauss. Thus, attention weights computed from these inner products do not depend on $R$ and generalize across different projections.
@@
$q_{i}=\tilde{\phi}_{i},\quad k_{i}=\tilde{\phi}_{i},\quad v_{i}=f_{v}(x_{i}).$
This ensures attention logits are based on inner products:
$e_{ij}=\frac{q^{\top}_{i}k_{j}}{\sqrt{d}}=\frac{\tilde{\phi}^{\top}_{i}\tilde{\phi}_{j}}{\sqrt{d}}=\frac{\phi^{\top}_{i}R^{\top}R\phi_{j}}{\sqrt{d}}\approx\frac{\phi^{\top}_{i}\phi_{j}}{\sqrt{d}},$
which are approximately gauge-invariant (see Fig. 1, left). By limiting the embeddings to the query-key computation and not using them as values, we ensure that downstream layers (which operate on node features) cannot access gauge-dependent information. Algorithm 1 in Section A.2 explains how we compute graph spectral positional embeddings and Algorithm 2 details the implementation.
@@
The Gauge-Invariant Spectral Self-Attention operation thus preserves gauge invariance, but at the cost of giving up a lot of the flexibility of regular self-attention. In particular, there is no mechanism that allows for a modification of the vectors $\tilde{\phi}_{i}$ through learning. In fact, applying even just a linear operation on $\tilde{\phi}_{i}$ would break again gauge invariance. However, notice that rescaling each $\tilde{\phi}_{i}$ by a scalar possibly depending on the node features $s(x_{i})\in\mathbb{R}$ would modify the similarity between graph positional embeddings in the same way across gauge choices, since scalars commute with orthogonal projections, meaning that it is an equivariant operation across gauges: $(s(x_{i})\tilde{\phi}_{i})^{\top}(s(x_{j})\tilde{\phi}_{j})=s(x_{i})s(x_{j})(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})=s(x_{i})s(x_{j})(\phi_{i}^{\top}\phi_{j})$.
@@
$q_{i}=f_{q}(x_{i}),\quad k_{i}=f_{k}(x_{i}),\quad v_{i}=\tilde{\phi}_{i},$
and the output in equation 1 such that it is constrained to operate on $\tilde{\phi}_{i}$, i.e. $\tilde{\phi}_{i}^{l+1}=\sum_{j=1}^{N}\alpha_{ij}v_{j}$, where $\tilde{\phi}_{i}^{l+1}$ indicates the graph positional encoding that will be used in the next layer $l+1$. Algorithm 3 in Section A.2 details how the implementation of Gauge-Equivariant Spectral Self-Attention relates to regular Self-Attention.
@@
Gauge-Invariant Spectral Self-Attention ensures that we can compute reliable graph positional encoding with linear time complexity in the number of nodes in the graph $N$. In order to maintain that linear scaling end-to-end, the very last component of our architecture aims to address the quadratic scaling of Transformers by implementing a linear version of self-attention. In particular, we implement the linear transformer by Katharopoulos2020.
Crucially, as feature map we use $\varphi(x)=\mathrm{ReLU}(x)$, which is a map that induces a kernel $k_{0}(\cdot)$ corresponding to the arc-cosine kernel (Cho2009). More specifically, for random features $\tilde{\phi}_{i},\tilde{\phi}_{j}\in\mathbb{R}^{r}$, the attention weights $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ converge to a kernel function that depends only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$. Since $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}\approx\phi_{i}^{\top}\phi_{j}$ by Johnson-Lindenstrauss (as established earlier), this preserves gauge invariance: attention weights depend only on gauge-invariant inner products between true spectral embeddings. For further considerations on the choice of the feature map $\varphi(\cdot)$ see the note in Appendix A.2.
In order to fully exploit the capabilities of linear attention and mitigate its drawbacks like the reported lack of sharp attention scores compared to softmax attention, we design a parallel architecture inspired from EfficientViT by Cai2022 who proposed a multi-scale linear attention architecture. Just like EfficientViT our Multi-Scale Gauge-Invariant Spectral Transformer Block has 3 parallel branches: a feature branch consisting in a linear transformer block acting on node features $x$ alone, a local branch consisting in a graph-convolution layer also acting on $x$ followed by a linear transformer block, and a global branch consisting in our Gauge-Invariant Spectral Self-Attention layer followed by a Gauge-Equivariant Spectral Self-Attention layer (which as explained act on both node features $x$ and graph positional embeddings $\tilde{\phi}$) then followed by a linear transformer. In keeping with the analogy with EfficientViT, the role of the graph-convolution layer (which simply averages node features across adjacent nodes) is to emphasize local information, which would be otherwise diffused by linear attention. Conversely, Gauge-Invariant Spectral attention integrates global information across the graph. This block is represented in the right panel of Fig. 1 and represents a unit layer that is sequentially replicated multiple times.
In Appendix A.5 we report ablation studies empirically showing that all branches meaningfully contribute to the final accuracy of the architecture, specifically for PPI where the full architecture achieves SOTA performance but would not if any of the branches were missing.

### 10 Apr 2026 v2

@@ Abstract
Adapting transformer positional encodings to graphs and meshes faces a fundamental tension: exact spectral methods require cubic-complexity eigendecomposition and inadvertently break gauge invariance through numerical solver artifacts, while existing efficient approximations sacrifice gauge symmetry by design. Both failure modes cause catastrophic generalization loss in inductive settings, where models fail when encountering different spectral decompositions of similar graphs or different discretizations of the same domain. We propose GIST (Gauge-Invariant Spectral Transformer), a graph transformer resolving this tension by restricting attention to pairwise inner products of efficient approximate spectral embeddings. We prove these inner products estimate an exactly gauge-invariant graph kernel at end-to-end $\mathcal{O}(N)$ complexity, and establish a formal connection between gauge invariance and discretization invariance: gauge invariance guarantees discretization-invariant learning with bounded mismatch error, making GIST the first scalable graph transformer with provable neural operator guarantees. Empirically, GIST matches state-of-the-art on standard graph benchmarks (e.g., achieving 99.50% micro-F1 on PPI) while uniquely scaling to mesh-based neural operator benchmarks with up to 750K nodes, achieving state-of-the-art on the AirfRANS, ShapeNet-Car, DrivAerNet, and DrivAerNet++ benchmarks.
@@ 1 Introduction
Following their incredible success for processing sequential data in Natural Language Processing, Transformers (Vaswani2017) have been demonstrating a remarkable capacity for handling data of increasing structural complexity. Lee2019 have proposed a variant of the transformer block for permutation invariant data with their Set Transformer architecture; Dosovitskiy2021 have adapted the self-attention mechanism to 2D images with the very influential Vision Transformer architecture; and Bertasius2021 have extended transformers to video analysis with their Video Vision Transformer, demonstrating how attention mechanisms capture both spatial and temporal dependencies across video frames. This progression from sequential text to increasingly structured data suggests that Transformers are poised to tackle even more complex data structures, including irregular meshes and graphs.
Indeed, recent developments in adapting Transformers to graphs have shown promising results in capturing global relationships through self-attention, overcoming the long-range dependency limitations of GNNs that rely on localized message passing (Dwivedi2022; Zhu2023c).
@@
Spectral embeddings depend on arbitrary choices: eigenvector signs, basis orderings, and solver artifacts. A neural network trained on embeddings from one set of choices will learn features that do not transfer when applied with different choices or to different graphs, fundamentally breaking generalization in inductive learning. This gauge dependence is particularly problematic when moving between different graph representations or numerical solvers.
This problem becomes critical for neural operators, where the input domain (a physical system or geometric shape) is discretized into a computational mesh. Different mesh resolutions induce different graph Laplacians and spectral decompositions, each with distinct gauge choices. Models trained on one discretization cannot transfer to another because spectral embeddings at different resolutions are incomparable and the attention kernels cannot be meaningfully compared across discretizations. Gauge invariance, making representations and learned operators independent of arbitrary spectral choices, is essential for both robust inductive learning and for constructing neural operators with provably bounded discretization error.
@@
We propose Gauge-Invariant Spectral Transformers (GIST), which overcomes both barriers through a key insight: gauge invariance can be preserved by restricting attention to inner products between spectral embeddings, which remain invariant even under approximate spectral computations. We show that a broad class of efficient $\mathcal{O}(N)$ spectral embeddings of the form $\widetilde{\Phi}=f(P)\cdot R$, where $f$ is any spectral filter and $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$, share a key property: their inner products are unbiased estimators of the graph kernel $K=f(P)f(P)^{\top}$, which depends only on the graph operator $P$ and is therefore exactly gauge-invariant. By restricting attention operations to these inner products and combining with linear attention, we achieve end-to-end $\mathcal{O}(N)$ complexity while recovering gauge invariance exactly at the population level, with statistical error $\mathcal{O}(r^{-1/2})$. Unlike prior approaches that address these barriers separately, GIST achieves gauge invariance and linear complexity simultaneously. Our contributions in short are:
@@
    identifying gauge invariance breaking as a failure mode shared by both exact and approximate spectral methods, and demonstrating that it causes catastrophic generalization loss in inductive graph settings;
@@
    proving that gauge invariance is necessary for discretization invariance with bounded discretization error and establishing the first discretization mismatch bound for a graph neural operator;
@@
    proposing GIST, the first graph transformer that simultaneously achieves gauge invariance, end-to-end $\mathcal{O}(N)$ complexity, and global spectral expressivity, enabling scalable neural operators with provable error bounds validated by state-of-the-art results on mesh problems with up to 750K nodes.
@@ 2 Related Works
Graphormer (Ying2021) introduces the idea of integrating structural encodings such as shortest path distances and centrality in Transformers. Similarly, Dwivedi2022 propose LSPE (Learnable Structural and Positional Encodings), an architecture that decouples structural and positional representations. Kreuzer2021 propose Spectral Attention Network (SAN), which introduces learned positional encodings from the full Laplacian spectrum. Park2022a develop Graph Relative Positional Encoding (GRPE), which extends relative positional encoding to graphs by considering features representing node-topology and node-edge interactions. Hierarchical Graph Transformer (Zhu2023c) addresses scalability to million-node graphs through graph hierarchies and coarsening techniques. SpecFormer (bo2023specformer) and PolyFormer (Chen2025) are recent spectral Transformers that improve over SAN by leveraging approximate spectral bases or low-rank polynomial Laplacian filters to enhance scalability and accuracy on graph tasks.
Several recent architectures aim to bridge local and global reasoning via structural encodings and scalable attention. GraphGPS (rampasek2022GPS) adopts a hybrid paradigm that decouples local message passing from global attention while integrating spectral positional encodings to capture multi-scale interactions. However, unlike GIST, it lacks a mechanism to algorithmically preserve gauge invariance, leaving the model susceptible to the arbitrary numerical artifacts and sign flips inherent in spectral decompositions. While models like Exphormer (shirzad2023exphormer) attempt to address the resulting complexity by replacing full attention with sparse expander-based mechanisms, they still struggle with the high overhead of global coordinate systems. Similarly, tokenized approaches such as TokenGT (hamilton2017inductive) and NAGphormer (chen2023nagphormer) treat graphs as sets of tokens with [CLS]-style readouts, yet these models scale poorly due to memory-intensive tokenization and positional encoding costs that grow super-linearly with graph size. Consequently, these frameworks are rarely evaluated on large-scale inductive tasks like ogbn-arxiv, where maintaining structural encodings becomes computationally prohibitive.
@@
Neural operators learn maps between continuous function spaces rather than fixed-dimensional vectors. Two properties are crucial here: (i) discretization invariance, i.e., a single set of parameters applies across discretizations (meshes, resolutions, and sampling locations) of the same underlying continuum problem; and (ii) global integration, i.e., the ability to represent nonlocal interactions via learned integral kernels, rather than being limited to finite-receptive-fields. Formally, a neural operator composes learned integral operators with pointwise nonlinearities, yielding universal approximation results for continuous nonlinear operators and implementations that share weights across resolutions. Our approach preserves these neural operator properties and improves scalability, allowing it to be applied to these cases (kovachki2023neuraloperator).
@@ 3.1 Preliminaries
The (normalized) graph Laplacian $\mathcal{L}=\mathds{1}-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$}} induces a natural metric via the resistance distance: $\Omega(i,j)=(e_{i}-e_{j})^{\top}\mathcal{L}^{\dagger}(e_{i}-e_{j})$, where $e_{i}$ is the $i$th standard basis vector and $\mathcal{L}^{\dagger}$ is the Moore-Penrose pseudoinverse (Klein1993). The resistance distance can be expressed via Laplacian eigenmaps, which satisfy:
@@
with $\lambda_{k},u_{k}$ being the eigenvalues and eigenvectors of $\mathcal{L}$. These eigenmaps are natural graph positional encodings (Dwivedi2021), but exact computation requires $\mathcal{O}(N^{3})$ eigendecomposition, prohibitive for large graphs.
@@
Note from equation 2 that the graph metric is determined by pairwise inner products of eigenmaps, since $\Omega(i,j)=\langle\phi_{i},\phi_{i}\rangle-2\langle\phi_{i},\phi_{j}\rangle+\langle\phi_{j},\phi_{j}\rangle$. This means we do not need the eigenvectors themselves, only an efficient way to estimate this kernel. A broad class of randomized methods achieves this, computing embeddings as
$\widetilde{\Phi}=f(P)\cdot R\;\in\mathbb{R}^{N\times r},$ (3)
where $P$ is a graph operator (e.g., the random-walk matrix $D^{-1}A$ or the normalized Laplacian) with eigendecomposition $P=U\Lambda U^{\top}$, $f$ is a spectral filter ($f(P)=U\,\mathrm{diag}(f(\lambda_{1}),\ldots,f(\lambda_{N}))\,U^{\top}$), $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$, and $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ is the $i$-th row of $\widetilde{\Phi}$. Laplacian eigenmaps equation 2 are the special case $f(\lambda)=\lambda^{-1/2}$, $P=\mathcal{L}$; polynomial filters, rational filters, and random-walk diffusions also fit this template, with $f(P)\cdot R$ computable via repeated sparse matrix-vector products without forming $f(P)$ explicitly.
A key property shared by all methods in this class is that their pairwise inner products are unbiased estimators of the graph kernel $K(i,j):=[f(P)f(P)^{\top}]_{ij}$:
$\mathbb{E}\!\left[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle\right]=K(i,j)=[f(P)f(P)^{\top}]_{ij},$ (4)
where the expectation is over the draw of $R$ and follows from $\mathbb{E}[RR^{\top}]=\mathds{1}$; when $P$ is symmetric this reduces to $K(i,j)=\sum_{l}f(\lambda_{l})^{2}\,(u_{l})_{i}(u_{l})_{j}$. Crucially, $K(i,j)$ is a function of $P$ alone and therefore exactly gauge-invariant: it is unchanged by sign flips, eigenspace rotations, and numerical solver choices (Proposition A.5). The estimation error is statistical and vanishes as $\mathcal{O}(1/\sqrt{r})$ (Lemma A.6). In our implementation we use FastRP (Chen2019e), a polynomial filter $f(P)=\sum_{i=1}^{k}P^{i}$ computed via sparse random projections and power iterations at $\mathcal{O}(Nrk)$ cost (see Algorithm 1), but any method in this class works equally well, e.g. graph diffusion kernels (Kondor2002), Chebyshev polynomial filters (Hammond2011), or randomized SVD-based embeddings (Halko2011).
@@
While the inner products of approximate spectral embeddings estimate a gauge-invariant kernel, the embeddings themselves depend on arbitrary choices: the random matrix $R$, eigensolver conventions, eigenvalue ordering. Any model operating on these embeddings directly, rather than on their inner products, will learn features that don’t transfer across graphs, resulting in a generalization failure. Our approach addresses this by using approximate spectral embeddings as positional encodings, but restricting the Transformer to operate only on their gauge-invariant inner products.
@@
The key insight is that spectral embeddings of the form $\widetilde{\Phi}=f(P)\cdot R$ produce inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ that are unbiased estimators of the exactly gauge-invariant kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ (equation 4). By restricting all operations to depend on the embeddings only through these inner products, gauge invariance is preserved exactly at the population level (with only statistical noise of order $\mathcal{O}(1/\sqrt{r})$) while maintaining computational efficiency.
Figure 1: Gauge-Invariant Spectral Transformer. Left: Gauge-Invariant Spectral Self-Attention operates on graph positional embeddings $\tilde{\phi}$ as queries and keys, and node features $x$ as values. The output of the self-attention operation is then combined with $x$ through a residual connection. Limiting $\tilde{\phi}$ to queries and keys preserves gauge invariance across the self-attention block. The diagram uses standard softmax attention for clarity of exposition; the actual GIST implementation replaces it with linear attention. Right: Gauge-Invariant Spectral Self-Attention is embedded in a Multi-Scale Gauge-Invariant Spectral Transformer Block which comprises 3 parallel branches inspired by EfficientViT.
@@
We now introduce our main contribution: Gauge-Invariant Spectral Transformer (GIST). The first ingredient of GIST is Gauge-Invariant Spectral Self-Attention, which operates on approximate spectral embeddings $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ (the $i$-th row of $\widetilde{\Phi}=f(P)\cdot R$) but restricts attention computations to inner products between embeddings. The key observation is that while the individual embeddings $\tilde{\phi}_{i}$ depend on the random matrix $R$, their inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ are unbiased estimators of the gauge-invariant kernel $K(i,j)$ (equation 4), which depends only on $P$. Attention weights computed from these inner products are therefore gauge-invariant and transfer across graphs.
@@
$q_{i}=\tilde{\phi}_{i},\quad k_{i}=\tilde{\phi}_{i},\quad v_{i}=W_{v}x_{i},$
where $W_{v}$ is a learned linear projection of the node features $x_{i}$, as usual in transformers. This ensures attention logits are based on inner products:
$e_{ij}=q^{\top}_{i}k_{j}=\tilde{\phi}^{\top}_{i}\tilde{\phi}_{j}=K(i,j)+\mathcal{O}(1/\sqrt{r}),$
where $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is the exactly gauge-invariant kernel from equation 4 (see Fig. 1, left). By limiting the embeddings to the query-key computation and not using them as values, we ensure that downstream layers (which operate on node features) cannot access gauge-dependent information. Algorithm 1 in Section B gives our specific implementation of the spectral embedding computation (which uses FastRP), and Algorithm 2 details the attention implementation. Notice that the softmax in equation 1 is replaced by linear attention in our implementation (see the Linear Self-Attention paragraph below).
@@
The Gauge-Invariant Spectral Self-Attention operation thus preserves gauge invariance, but at the cost of giving up a lot of the flexibility of regular self-attention. In particular, there is no mechanism that allows for a modification of the vectors $\tilde{\phi}_{i}$ through learning. In fact, applying even just a linear operation on $\tilde{\phi}_{i}$ would break again gauge invariance. However, notice that rescaling each $\tilde{\phi}_{i}$ by a scalar possibly depending on the node features $s(x_{i})\in\mathbb{R}$ would modify the similarity between graph positional embeddings in the same way across gauge choices, since scalars commute with orthogonal projections, meaning that it is an equivariant operation across gauges: $(s(x_{i})\tilde{\phi}_{i})^{\top}(s(x_{j})\tilde{\phi}_{j})=s(x_{i})s(x_{j})(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})=s(x_{i})s(x_{j})\cdot K(i,j)+\mathcal{O}(1/\sqrt{r})$.
@@
$q_{i}=W_{q}x_{i},\quad k_{i}=W_{k}x_{i},\quad v_{i}=\tilde{\phi}_{i},$
and the output in equation 1 such that it is constrained to operate on $\tilde{\phi}_{i}$, i.e. $\tilde{\phi}_{i}^{l+1}=\sum_{j=1}^{N}\alpha_{ij}v_{j}$, where $\tilde{\phi}_{i}^{l+1}$ indicates the graph positional encoding that will be used in the next layer $l+1$. Algorithm 3 in Section B details how the implementation of Gauge-Equivariant Spectral Self-Attention relates to regular Self-Attention.
@@
Gauge-Invariant Spectral Self-Attention ensures that we can compute reliable graph positional encoding with linear time complexity in the number of nodes in the graph $N$. In order to maintain that linear scaling end-to-end, all GIST attention blocks replace softmax attention (introduced above and depicted in Fig. 1, left, for clarity) with the linear transformer by Katharopoulos2020, using a feature map $\varphi$ so that $\alpha_{ij}\propto\langle\varphi(q_{i}),\varphi(k_{j})\rangle$.
Crucially, as feature map we use $\varphi(x)=\mathrm{ReLU}(x)$, which is a map that induces a kernel $k_{0}(\cdot)$ corresponding to the arc-cosine kernel (Cho2009). More specifically, for random features $\tilde{\phi}_{i},\tilde{\phi}_{j}\in\mathbb{R}^{r}$, the attention weights $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ converge to a kernel function that depends only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$. Since $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$ estimates the exactly gauge-invariant kernel $K(i,j)$ (equation 4), this preserves gauge invariance: attention weights depend only on gauge-invariant quantities. For further considerations on the choice of the feature map $\varphi(\cdot)$ see the note in Appendix B.
To mitigate linear attention’s drawbacks (notably lack of sharp attention scores compared to softmax attention), we design a parallel architecture inspired by EfficientViT (Cai2022), a multi-scale linear attention architecture. Our Multi-Scale Gauge-Invariant Spectral Transformer Block comprises three parallel branches: a feature branch (linear transformer on node features $x$), a local branch (graph convolution on $x$ followed by linear transformer), and a global branch (Gauge-Invariant Spectral Self-Attention followed by Gauge-Equivariant Spectral Self-Attention operating on both $x$ and $\tilde{\phi}$, then linear transformer). The local branch emphasizes local information that linear attention would otherwise diffuse, while the global branch integrates global graph information. See Figure 1 (right panel) for the architecture. Ablation studies in Appendix C.3 confirm that each branch meaningfully contributes to performance.